please answer the following questions
Answer #1

------------------------

so order is , 276.08 nm , 497 nm , 828.25 nm

apply V = Lf

so wavelength L = 3e8/800 e6

L = 0.375 m

apply d sin theta = mL

sin theta = 0.375/13

sin theta = 0.0288

theta = 1.652 deg
----------------------------------------

apply Y = 1.22 LR/D


R = YD/1.22L

here Y = 0.8/2 = 0.4 cm ( fron centre to one end)

R = 0.4 cm *0.15 mm /(1.22* 633nm)

R = 0.776m or 77.6 cm